mary is 5 times as old as her daughter, and. three years ago she was 9 times as old as ann. How old is mary
Answer by jorel1380(3719)

Let a be ann. Then mary is 5a. So:
5a-3=9(a-3)
5a-3=9a-27
4a=24
a=6
ann is 6
mary is 30. ☺☺☺☺